OpenStudy (anonymous):

Bob has 19 Pokemon cards. Suzy has an unknown number. Joe has 1/4 of the number that Bob+Suzy have. There are 35 cards altogether. How many does Suzy have? I need Help my friend gave me this question and I don't know how to solve It Please Help... I'll give a Medal and Fan you!! <3

OpenStudy (anonymous):

And also can someone show me how to solve it step by step?? <3

OpenStudy (anonymous):

Pokemon cards?!

OpenStudy (anonymous):

the way id do it...make an equation for the sentence "Joe has 1/4 of the number that Bob+Suzy have." call joe j and suzy x 1/4(19+x)=j and in total they have 35 cards so 19+j+x=35 now we have two equations

OpenStudy (anonymous):

j is solved for in the first equation so we can take that and plug it into the second one for j 19+1/4(19+x)+x=35 now can you solve for x?

OpenStudy (anonymous):

yes I can solve :)

OpenStudy (anonymous):

is 19 + 1/4 = 19 1/4? @CalebBeavers

OpenStudy (anonymous):

okay I don't get it?

OpenStudy (anonymous):

The 1/4 is being multiplied, distribute it into the parentheses

OpenStudy (anonymous):

so 19/1 * 1/4

OpenStudy (anonymous):

4 3/4?

OpenStudy (anonymous):

Multiply straight across 19/4 and also multiply it by the x so we have 19+(19/4)+(1/4)x+x=35

OpenStudy (anonymous):

Now just combine like terms and isolate x 19+(19/4)+(5/4)x=35 95/4+(5/4)x=35 (5/4)x=45/4 x=(45/4)*(4/5)

OpenStudy (kirbykirby):

/Ok Let B = Bob, J = Joe and S = Suzy B = 19 S = x J = (1/4) (x+19) Together they have 35: 19+x+(1/4)(x+19) = 35 19 + x + x/4 + 19/4 = 35 and solve for x :) : 5x/4 + 95/4 = 35 5x/4 = 35 - 95/4 5x/4 = 45/4 5x = 45 x = 9 Then just substitute the value for x for each person: we get that x= 9 So S = x = 9